OpenStudy (anonymous):

How do you solve this: log(x-1)log(x-10)+log(x-1)-2log(x-10)=2 I tried and didn't succeed.

OpenStudy (asnaseer):

if you let a=log(x-1) and b=log(x-10) then you can rewrite your equation as: ab+a-2b=2 which can be manipulated to become: ab +a - 2b - 2 = 0 see if you can factorise this